write an equation of the line iwth the given slope and the given y-intercept. Leave the answer in slope-intercept form.
slope -2/3, y-intercept (0, 5)

Answer by mananth(16946)

m= - 2/ 3

Plug value of the slope and point ( 0 , 5 ) in
Y = m x + b
5.00 = -0 + b
b= 5 - 0
b= 5
So the equation will be
Y = - 2/3 x + 5
m.ananth@hotmail.ca